OpenStudy (anonymous):

Angie adds 7 to the product of 8 and a number, and gets 39 as the result. Which expression can be used to determine the value of the number?

OpenStudy (anonymous):

A. 7 + 8 + n = 39 B. n + 7 × 8 = 39 C. 7 + 8n = 39 D. 7n + 8 = 39

OpenStudy (anonymous):

help